# Finite-Time Analysis of Gradient Descent for Shallow Transformers

### v1

@@ Abstract
Understanding why Transformers perform so well remains challenging due to their non-convex optimization landscape. In this work, we analyze a shallow Transformer with $m$ independent heads trained by projected gradient descent in the kernel regime. Our analysis reveals two main findings: (i) the width required for nonasymptotic guarantees scales only logarithmically with the sample size $n$, and (ii) the optimization error is independent of the sequence length $T$. This contrasts sharply with recurrent architectures, where the optimization error can grow exponentially with $T$. The trade-off is memory: to keep the full context, the Transformer’s memory requirement grows with the sequence length. We validate our theoretical results numerically in a teacher–student setting and confirm the predicted scaling laws for Transformers.
@@ 1 INTRODUCTION
Transformers have reshaped modern machine learning, achieving state-of-the-art results in language, vision, and multi-modal tasks (Vaswani et al., 2017; Devlin et al., 2019; Brown et al., 2020; Dosovitskiy et al., 2021). Despite this empirical success, a sharp nonasymptotic understanding of their training dynamics remains elusive. Two obstacles stand out: (i) the highly nonlinear nature of the self-attention layer, whereby each output coordinate depends on all inputs through a softmax reweighting, making standard arguments for fully-connected networks or CNNs difficult to apply; and (ii) the nonconvexity of the training landscape, which complicates guarantees for gradient-based methods outside of infinite-width limits.
@@
where $\sigma_{s}$ denotes the column-wise softmax, and $\beta$ is the softmax scaling. In the original Transformer, one takes $\beta=m^{-1/2}$ to stabilize the dot-product variance (Vaswani et al., 2017). However, to make the analysis of (1) tractable, prior work often studies variants that take one of the following three directions: (i) avoid $\sigma_{s}$ entirely to obtain linear attention (Katharopoulos et al., 2020; Choromanski et al., 2021; Lu et al., 2021; Zheng et al., 2022; Zhang et al., 2024); (ii) replace $\sigma_{s}$ with a pointwise activation such as ReLU (Wortsman et al., 2023; Hron et al., 2020); or (iii) take infinite-width limits under $\beta=m^{-1}$, in which the softmax weights approach uniform distribution and the attention layer reduces to mean-pooling (Wu et al., 2023; Noci et al., 2023; Bordelon et al., 2024). While analytically convenient, these modifications limit the scope of the resulting guarantees.
@@
More recently, an interesting work by Wu et al. (2023) established global convergence for shallow Transformers at finite width under the realistic $\beta=m^{-1/2}$ scaling, showing that cubic overparameterization in the sample size $n$ suffices for gradient descent to converge at a linear rate. Their proof follows the usual neural tangent kernel (NTK) recipe of enforcing a strict positive definiteness to a Gram matrix to drive convergence (Du et al., 2019a, b; Allen-Zhu et al., 2019; Zou and Gu, 2019; Nguyen, 2021; Nguyen et al., 2021). In particular, Wu et al. (2023) control the minimum singular value of a feature matrix built from the last hidden layer, which relies on the regime $d\geq n$, thereby limiting its practicality.
@@
Parallel to these developments for Transformers, new techniques have sharpened finite-time analyses for recurrent architectures (Cayci and Eryilmaz, 2024a, b). Cayci and Eryilmaz (2024a) prove nonasymptotic convergence of gradient descent for IndRNNs with dramatically improved width requirements—most notably, width scaling only logarithmically in the sample size $n$. Their proof is organized around norm-constrained transportation mappings that represent the target dynamical system in the RKHS induced by the RNN’s NTK, tying random-feature ideas to near-initialization training (Rahimi and Recht, 2007; Ji et al., 2020). Together with a projection step that keeps the gradient-descent iterates near initialization, they obtain finite-sample, finite-width convergence. The bounds also make explicit the price of long-term dependencies: beyond a cutoff governed by the activation’s Lipschitz constant, both iteration complexity and the required width can grow exponentially with the sequence length $T$.
@@
In this paper, building on techniques from Cayci and Eryilmaz (2024a), we analyze a shallow multi-head Transformer trained by projected gradient descent algorithms (ProjGD and ProjSGD) in the near-initialization regime while preserving the nonlinearity of the attention layer. We address the nonlinearity and nonconvexity of the problem by working with projected updates that keep parameters in a small neighborhood of initialization. In this regime, a transportation-mapping viewpoint links the random features model induced at initialization to the NTK linearization and provides uniform control of the linearization error. We then exploit an explicit gradient characterization of the attention pathway to obtain optimization bounds independent of sequence length $T$. The main contributions of this work can be highlighted as follows:
@@
•  Bypassing positive definiteness of the NTK. Unlike the original works in the NTK regime, our analysis does not require strict positive definiteness of the NTK.
@@
•  Optimization error independent of input sequence length $T$. Our ProjGD bound does not degrade with the sequence length $T$. This contrasts sharply with RNN analyses, where the dependence on $T$ can be exponential under long-term dependencies. The trade-off is memory: retaining full context implies that the Transformer’s memory footprint grows with $T$ due to attention over all tokens.
@@ 2 RELATED WORK
Representation‑theoretic results establish that Transformers are universal approximators of permutation‑equivariant sequence‑to‑sequence maps, and, with trainable positional encodings, of general continuous sequence‑to‑sequence functions on compact domains (Yun et al., 2020). At infinite width, attention admits NTK descriptions under suitable scalings: Hron et al. (2020) derive attention NTK limits, showing that under the common $\beta=m^{-1/2}$, letting the number of heads grow results in a deterministic NTK. Under $\beta=m^{-1}$, deterministic kernels can exist but may degenerate toward mean-pooling or require $W_{Q}=W_{K}$ (Yang, 2020). Wu et al. (2023) complement their finite‑width convergence proof with an NTK analysis for shallow Transformers with $\beta=m^{-1}$.
@@
Wu et al. (2023) prove global convergence of a shallow Transformer while preserving the softmax attention nonlinearity and carefully diagnosing scaling/initialization choices. Complementary results capture in‑context learning dynamics with true softmax: Huang et al. (2024) give finite‑time guarantees and a phase‑wise view of attention; Song et al. (2024) show convergence under suitable initialization and reveal that softmax can introduce spurious local minima whereas a Gaussian attention smooths the landscape. Simplified settings using linear attention establish convergence for linear predictors learned in‑context (Zhang et al., 2024). Beyond linear tasks, Shen et al. (2024) obtain linear‑rate convergence to the Bayes classifier for in‑context classification with a shallow multi‑head Transformer. Further optimization views sharpen what attention learns: Tarzanagh et al. (2023) show that gradient descent on a single attention layer aligns with a max‑margin (SVM‑like) token‑selection rule. Nichani et al. (2024) analyze a two‑layer attention‑only Transformer and prove that gradient descent recovers the latent causal structure of the in-context learning problem.
@@
For $k\in\mathbb{N}$ we write $k]=\{1,2,\dots,k\}$. For matrices $A,B$ we write $\langle A,B\rangle_{F}=\mathrm{tr}(A^{\top}B)$ and $\|A\|_{F}$ for the Frobenius norm. The probability simplex in $\mathbb{R}^{T}$ is $\Delta_{T}=\{\,\alpha\in\mathbb{R}^{T}:\alpha_{t}\geq 0,\ \sum_{t=1}^{T}\alpha_{t}=1\,\}$ and $\mathbb{R}_{+}=(0,\infty)$ is the set of positive reals. For a vector $x$, $\operatorname{diag}(x)$ denotes the diagonal matrix with diagonal $x$. For matrices $A_{1},\ldots,A_{r}$, $\operatorname{blkdiag}(A_{1},\ldots,A_{r})$ denotes the block-diagonal matrix with diagonal blocks $A_{1},\ldots,A_{r}$ and zero elsewhere. For $A\in\mathbb{R}^{m\times n}$, $\operatorname{vec}(A)\in\mathbb{R}^{mn}$ denotes the column-wise vectorization of $A$. The notation $\lesssim_{a,b}$ omits the constants depending on the parameters $a,b$. For vectors $x,y$, $x\leq y$ denotes element-wise comparison. $\mathrm{Unif}(V)$ denotes the uniform distribution over the set $V$, and $\mathrm{Rad}(\alpha)$ denotes $\mathrm{Unif}(\{-\alpha,\alpha\})$.
@@ 3 PROBLEM SETUP
We specify the objective, introduce the shallow multi-head Transformer studied in this work, and formalize the supervised learning task with the optimization algorithms that will be analyzed.
@@ 3.2 Transformer Architecture
Fix $X\in\mathcal{X}$. Since each training sample has a scalar target, we employ pooled self-attention driven by a single query $q_{X}\in\mathbb{R}^{d}$ with $\|q_{X}\|_{2}\leq 1$. The query $q_{X}$ is not trainable and may be a fixed vector ([CLS] as in BERT (Devlin et al., 2019), or ViT’s class token (Dosovitskiy et al., 2021)), a deterministic function of the input (e.g., last-token pooling $q_{X}=X_{T}$ for next-token prediction (Radford et al., 2019)), or a pre-specified seed query used to summarize high-dimensional inputs (Jaegle et al., 2021).
For $i\in[m]$ and trainable $W_{i}\!\in\!\mathbb{R}^{d\times d}$, $U_{i}\!\in\!\mathbb{R}^{d}$, and $c_{i}\!\in\!\mathbb{R}$, we consider the following model.
@@
where $\sigma_{s}:\mathbb{R}^{T}\!\to\!\Delta_{T}$ is the softmax:
$(\sigma_{s}(z))_{t}\;=\;\frac{\exp(z_{t})}{\sum_{s=1}^{T}\exp(z_{s})}.$
@@
$f(X;\varphi)\coloneqq\tfrac{1}{\sqrt{m}}\sum_{i=1}^{m}c_{i}\,h(X;\theta_{i})\in\mathbb{R},$ (4)
@@
Following prior analyses (Wu et al., 2023; Tarzanagh et al., 2023; Nichani et al., 2024; Oymak et al., 2023) we combine the query and key matrices in (1) into a single attention weight matrix $W_{i}$, which corresponds to the canonical $\beta=m^{-1/2}$ scaling, and absorb the value matrix into the next hidden layer’s weights $U_{i}$.
@@
In standard multi-head attention one concatenates $h$ attention heads into $A(X)=(a(X;W_{1}),\ldots,a(X;W_{h}))\in\mathbb{R}^{hd}$ and applies a feed-forward layer with a weight matrix $V\in\mathbb{R}^{m\times hd}$, resulting in $\sigma(VA(X))$. Our architecture is equivalent to this with a block-diagonal weight matrix with one neuron per head, i.e., $h=m$ and $V=\operatorname{blkdiag}(U_{1}^{\top},\ldots,U_{m}^{\top})$. This independent head constraint acts as a structural regularizer: it preserves per-head interpretability, reduces cross-head interference, restores a deterministic NTK, and simplifies the analysis without requiring any of the architectural changes mentioned in Section 1.
Remark 3.
A key point to establish our main results will be to show the local Lipschitz continuity and smoothness of the mapping $\theta_{i}\mapsto h(X;\theta_{i})$ in Lemma 2.
@@ 3.3 Algorithms for Empirical Risk Minimization
We analyze (i) ProjGD with $B_{s}=[n]$ for all $s$, and (ii) ProjSGD with $B_{s}=\{J_{s}\}$ with $J_{s}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathrm{Unif}([n])$.

### v2

@@ Abstract
Understanding why Transformers perform so well remains challenging due to their non-convex optimization landscape. In this work, we analyze a shallow Transformer with $m$ independent heads trained by projected gradient descent in the kernel regime. Our analysis reveals two main findings: (i) the width required for nonasymptotic guarantees scales only logarithmically with the sample size $n$, and (ii) the optimization error is independent of the sequence length $T$. This contrasts sharply with recurrent architectures, where the optimization error can grow exponentially with $T$. The trade-off is memory: to keep the full context, the Transformer’s memory requirement grows with the sequence length. We validate our theoretical results numerically in a teacher–student setting and compare Transformers with recurrent architectures on an autoregressive task.
@@ 1 INTRODUCTION
Transformers have reshaped modern machine learning, achieving state-of-the-art results in language, vision, and multimodal tasks (Vaswani et al., 2017; Devlin et al., 2019; Brown et al., 2020; Dosovitskiy et al., 2021). Despite this empirical success, a sharp nonasymptotic understanding of their training dynamics remains elusive. Two obstacles stand out: (i) self-attention is highly nonlinear—each output coordinate depends on all inputs through a softmax reweighting—so standard arguments for fully-connected networks or CNNs do not apply directly; and (ii) the training landscape is nonconvex, complicating guarantees for gradient-based methods beyond infinite-width limits.
@@
where $\sigma_{s}$ denotes the column-wise softmax, and $\beta$ is the softmax scaling. In the original Transformer, one takes $\beta=m^{-1/2}$ to stabilize the dot-product variance (Vaswani et al., 2017). However, to make the analysis of (1) tractable, prior work often studies variants that take one of the following directions: (i) avoid $\sigma_{s}$ entirely to obtain linear attention (Katharopoulos et al., 2020; Choromanski et al., 2021; Lu et al., 2021; Zheng et al., 2022; Zhang et al., 2024); (ii) replace $\sigma_{s}$ with a pointwise activation such as ReLU (Wortsman et al., 2023; Hron et al., 2020); or (iii) take infinite-width limits under $\beta=m^{-1}$, in which the softmax weights approach uniform distribution and the attention layer reduces to mean-pooling (Wu et al., 2023; Noci et al., 2023; Bordelon et al., 2024). While analytically convenient, these modifications limit the scope of the resulting guarantees.
@@
More recently, Wu et al. (2023) established global convergence for shallow Transformers at finite width under the realistic $\beta=m^{-1/2}$ scaling, showing that quadratic overparameterization in the sample size $n$ suffices for gradient descent to converge at a linear rate. Their proof follows the standard recipe of enforcing strict positive definiteness of a Gram matrix to drive convergence (Du et al., 2019a, b; Allen-Zhu et al., 2019; Zou and Gu, 2019; Nguyen, 2021; Nguyen et al., 2021). In particular, Wu et al. (2023) control the minimum singular value of a feature matrix built from the last hidden layer, which relies on the regime $d\geq n$, thereby limiting its practicality.
@@
Parallel to these developments for Transformers, recent work has sharpened finite-time analyses for recurrent architectures (Cayci and Eryilmaz, 2024a, b). In particular, Cayci and Eryilmaz (2024a) prove nonasymptotic convergence of gradient descent for independently recurrent neural networks (IndRNNs, Li et al., 2018) with width scaling only logarithmically in the sample size $n$, using transportation mappings to represent targets in the associated kernel function space. Their bounds also quantify the cost of long-term dependencies: beyond an activation-controlled cutoff, both the required width and iteration complexity can grow exponentially with the sequence length $T$.
@@
In this paper, building on techniques from Cayci and Eryilmaz (2024a), we analyze a shallow multi-head Transformer trained by gradient descent in the near-initialization regime while preserving the nonlinearity of the attention layer. The main contributions of this work can be highlighted as follows:
@@
•  Bypassing positive definiteness of the NTK/Gram. Unlike the original works in the NTK regime, our analysis does not require strict positive definiteness of the NTK/Gram matrix.
@@
•  Optimization error independent of input sequence length $T$. Our convergence bound does not degrade with the sequence length $T$. This contrasts sharply with RNN analyses, where the dependence on $T$ can be exponential under long-term dependencies. The trade-off is memory: retaining full context implies that the Transformer’s memory footprint grows with $T$ due to attention over all tokens.
@@
Technical challenges and ideas.
(i) We preserve the genuine softmax attention nonlinearity while avoiding any uniformly positive definite Gram/NTK assumption by working in a near-initialization regime, where the linearization error can be controlled under an explicit width requirement (Lemma 3). (ii) Within this NTK viewpoint, transportation mappings (Section 5.4) provide a convenient representation of targets in the Transformer-NTK RKHS and yield an approximation bound with only logarithmic dependence on the sample size (Lemma 5). (iii) To obtain bounds independent of the sequence length $T$, we exploit the covariance structure induced by the softmax Jacobian to bound attention gradient/Hessian norms uniformly in $T$ (Remark 3); this feeds into $T$-independent local Lipschitz/smoothness constants (Lemma 2) and ultimately yields the $T$-independent optimization error (Theorem 1).
All proofs of technical results are deferred to the appendix, which also provides additional details.
@@ 2 RELATED WORK
From a representation-theoretic perspective, Transformers are universal approximators of permutation-equivariant sequence-to-sequence maps, and, when equipped with trainable positional encodings, of general continuous sequence-to-sequence functions on compact domains (Yun et al., 2020). At infinite width, attention admits NTK descriptions under suitable scalings: Hron et al. (2020) derive attention NTK limits, showing that under the common $\beta=m^{-1/2}$, letting the number of heads grow results in a deterministic NTK. Under $\beta=m^{-1}$, deterministic kernels can exist but may degenerate toward mean-pooling or require $W_{Q}=W_{K}$ (Yang, 2020). Wu et al. (2023) complement their finite‑width convergence proof with an NTK analysis for shallow Transformers with the degenerate scaling $\beta=m^{-1}$.
@@
Wu et al. (2023) prove global convergence of a shallow Transformer while preserving the softmax attention nonlinearity and carefully diagnosing scaling/initialization choices. Deora et al. (2024) derive finite-time training and generalization bounds for gradient descent on a single-layer multi-head self-attention model in binary classification, where they obtain polylogarithmic head requirements under realizability and initialization conditions. Complementary results capture in‑context learning dynamics with true softmax: Huang et al. (2024) give finite‑time guarantees and a phase‑wise view of attention; Song et al. (2024) show convergence under suitable initialization and reveal that softmax can introduce spurious local minima whereas a Gaussian attention smooths the landscape. Simplified settings using linear attention establish convergence for linear predictors learned in‑context (Zhang et al., 2024). Beyond linear tasks, Shen et al. (2024) obtain linear‑rate convergence to the Bayes classifier for in‑context classification with a shallow multi‑head Transformer. Further optimization views sharpen what attention learns: Tarzanagh et al. (2023) show that gradient descent on a single attention layer aligns with a max‑margin (SVM‑like) token‑selection rule. Vasudeva et al. (2025) study implicit bias and convergence rates for single-layer self-attention in binary classification. Nichani et al. (2024) analyze a two‑layer attention‑only Transformer and prove that gradient descent recovers the latent causal structure of the in-context learning problem.
Comparison to Wu et al. (2023).
Our work is closest in spirit to Wu et al. (2023), which also studies shallow Transformers with genuine softmax attention under a regression setting. The main differences are: (i) Regime: Wu et al. (2023) prove global convergence via a well-conditioned last-layer feature/Gram matrix, whereas we work in a projected near-initialization NTK regime and do not assume a positive definite Gram matrix. (ii) Rates/width: their result achieves a linear rate with polynomial overparameterization in $n$ (Wu et al., 2023, Proposition 1), while our bounds have the canonical $1/\sqrt{\tau}$ optimization rate and require only logarithmic dependence on $n$ in the width (Theorem 1). (iii) Sequence length: our optimization constants are explicitly independent of $T$, while their conditions involve $T$-dependent quantities. (iv) NTK features: they complement their analysis with an NTK study under $1/m$-type scaling, which yields a vanishing attention component (Wu et al., 2023, Lemma 1); whereas our NTK decomposition retains a nontrivial attention block (Lemma 4).
@@
For $k\in\mathbb{N}$ we write $k]=\{1,2,\dots,k\}$. For matrices $A,B$ we write $\langle A,B\rangle_{F}=\mathrm{tr}(A^{\top}B)$ and $\|A\|_{F}$ for the Frobenius norm. The probability simplex in $\mathbb{R}^{T}$ is $\Delta_{T}=\{\,\alpha\in\mathbb{R}^{T}:\alpha_{t}\geq 0,\ \sum_{t=1}^{T}\alpha_{t}=1\,\}$ and $\mathbb{R}_{+}=(0,\infty)$. For a vector $x$, $\operatorname{diag}(x)$ denotes the diagonal matrix with diagonal $x$. For matrices $A_{1},\ldots,A_{r}$, $\operatorname{blkdiag}(A_{1},\ldots,A_{r})$ denotes the block-diagonal matrix with diagonal blocks $A_{1},\ldots,A_{r}$ and zero elsewhere. For $A\in\mathbb{R}^{m\times n}$, $\operatorname{vec}(A)\in\mathbb{R}^{mn}$ denotes the column-wise vectorization of $A$. The notation $\lesssim_{a,b}$ omits the constants depending on the parameters $a,b$. For vectors $x,y$, $x\leq y$ denotes element-wise comparison. $\mathrm{Unif}(V)$ denotes the uniform distribution over the set $V$, and $\mathrm{Rad}(\alpha)$ denotes $\mathrm{Unif}(\{-\alpha,\alpha\})$.
@@ 3 PROBLEM SETUP
We start by specifying the objective, introduce the shallow multi-head Transformer studied in this work, and formalize the optimization algorithms analyzed throughout the paper.
@@ 3.2 Transformer Architecture
Fix $X\in\mathcal{X}$. Since each training sample has a scalar target, we employ pooled self-attention driven by a single query $q_{X}\in\mathbb{R}^{d}$ with $\|q_{X}\|_{2}\leq 1$. The query $q_{X}$ is not trainable: it may be a fixed vector ([CLS] as in BERT (Devlin et al., 2019), or ViT’s class token (Dosovitskiy et al., 2021)), a deterministic function of the input (e.g., last-token pooling $q_{X}=X_{T}$ for next-token prediction (Radford et al., 2019)), or a pre-specified seed query used to summarize high-dimensional inputs (Jaegle et al., 2021).
Let $m\in\mathbb{N}$ be the width of our Transformer. For $i\in[m]$ and trainable $W_{i}\!\in\!\mathbb{R}^{d\times d}$, $U_{i}\!\in\!\mathbb{R}^{d}$, $c_{i}\!\in\!\mathbb{R}$, we consider the following model.
@@
where $\sigma_{s}:\mathbb{R}^{T}\!\to\!\Delta_{T}$ is the softmax function: $(\sigma_{s}(z))_{t}=\tfrac{\exp(z_{t})}{\sum_{s=1}^{T}\exp(z_{s})}.$
@@
$f(X;\varphi)\coloneqq\frac{1}{\sqrt{m}}\sum_{i=1}^{m}c_{i}\,h(X;\theta_{i})\in\mathbb{R},$ (4)
@@
Following prior analyses (Wu et al., 2023; Deora et al., 2024; Nichani et al., 2024; Oymak et al., 2023), we simplify the standard $(W_{Q},W_{K},W_{V})$ parameterization in (1). For pooled attention with a single query, the usual logits can be written as $\beta\,(W_{K}X)^{\top}(W_{Q}q_{X})\in\mathbb{R}^{T}$ and hence $\beta\,(W_{K}X)^{\top}(W_{Q}q_{X})=X^{\top}(\beta W_{K}^{\top}W_{Q})q_{X}$. Thus, collapsing $W_{Q}$ and $W_{K}$ into a single matrix in (2) amounts to setting $W_{i}:=\beta W_{K,i}^{\top}W_{Q,i}$. By initializing $W_{i}$ with bounded-variance entries, the scaling $\beta=1/\sqrt{m}$ is absorbed into the initialization and does not appear explicitly in (2). In contrast, $1/m$ scaling would drive logits toward zero and make the softmax weights nearly uniform, degenerating attention toward mean pooling (cf. Section 1). Finally, in this one-hidden-layer setting we absorb the value/output matrices into the subsequent linear map $U_{i}$.
@@
In standard multi-head attention one concatenates $h$ attention heads into $A(X)=(a(X;W_{1}),\ldots,a(X;W_{h}))\in\mathbb{R}^{hd}$ and applies a feed-forward layer with a weight matrix $V\in\mathbb{R}^{m\times hd}$, resulting in $\sigma(VA(X))$. Our architecture is equivalent to this with a block-diagonal weight matrix with one neuron per head, i.e., $h=m$ and $V=\operatorname{blkdiag}(U_{1}^{\top},\ldots,U_{m}^{\top})$. This independent head constraint acts as a structural regularizer: it preserves per-head interpretability, reduces cross-head interference, and restores a deterministic NTK, thereby simplifying the analysis without requiring any of the architectural changes mentioned in Section 1.
@@ 3.3 Algorithms for Empirical Risk Minimization
where $\Pi_{\Omega}$ denotes projection onto the set $\Omega$. We analyze (i) ProjGD with $B_{s}=[n]$ for all $s$, and (ii) ProjSGD with $B_{s}=\{J_{s}\}$ with $J_{s}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathrm{Unif}([n])$.
Why projection?
Our analysis is local around the random initialization: all bounds are proved on the neighborhood $\Omega_{\rho}$ (Definition 2), where (i) the NTK linearization error is uniformly controlled (Lemma 3) and (ii) the change-of-feature terms in the optimization argument can be bounded using local Lipschitz/smoothness constants (Lemma 2). Projecting the iterates onto $\Omega_{\rho}$ is therefore a simple analytical device that enforces this near-initialization regime. In practice, other mechanisms that keep parameters close to initialization (e.g., weight decay, explicit norm regularization, or early stopping) can play a similar role, and in Section 6 we also report experiments trained without projection.